### v1

@@ Abstract
[Summary] Purpose: The long scan times of quantitative MRI techniques make motion artifacts more likely. For MR-Fingerprinting-like approaches, this problem can be addressed with self-navigated retrospective motion correction based on reconstructions in a singular value decomposition (SVD) subspace. However, the SVD promotes high signal intensity in all tissues, which limits the contrast between tissue types and ultimately reduces the accuracy of registration. The purpose of this paper is to rotate the subspace for maximum contrast between two types of tissue and improve the accuracy of motion estimates.
@@
Results: A comparative analysis shows that the contrast-optimized basis significantly improve the parenchyma-CSF contrast, leading to smoother motion estimates and reduced artifacts in the quantitative maps.
@@
articletype: Pre-print - Submitted to Magnetic Resonance in Medicine00footnotetext: * E. Marchetto and S. Flassback contributed equally to this work.
@@ 1 Introduction
Magnetic resonance fingerprinting (MRF) is a multiparametric quantitative magnetic resonance imaging (qMRI) approach. Its key concept is a variation of sequence parameters between TRs to maintain a transient state of the magnetization. 1, 2
Motion-induced artifacts pose severe challenges in MRF-like experiments, whether due to physiological motion or involuntary motion due to pathological conditions. 3D acquisitions mitigate the risk of through-plane motion3 - which is prohibitively difficult to correct retrospectively - but the associated long scan times make motion more likely.
Fat-navigators have previously been proposed for robust retrospective motion correction in structural brain MRI,4, 5, 6 and they have successfully been implemented in a 3D MRF sequence.7 However, navigator-based motion correction techniques depend on the sequence timing, as they are typically acquired during waiting periods. Moreover, even fat-selective excitation pulses can perturb the spin dynamics via the magnetization transfer effect, which can compromise quantitative measurements. An alternative approach is self-navigated motion correction, which requires no additional data as the motion parameter estimation is derived from the acquired data itself. In 3D MRF-flavored acquisitions, a self-navigated motion correction approach was previously proposed. 8 Kurzawski et al. reconstructed brain navigators from 7 s segments in a sub-space spanned by the truncated singular value decomposition (SVD) subspace. This method utilizes the low-rank nature of the underlying data and produces coefficient images which are used to extract the motion estimates using rigid registration. However, the SVD promotes high signal intensity in all types of tissue, which limits the contrast and can ultimately reduce the accuracy of the registration and extracted motion estimates.
In this work, we aim to improve the accuracy of the motion estimates by deriving a basis that maximizes the signal of the fingerprints corresponding to brain parenchyma (i.e., white and gray matter), while minimizing the signal of fingerprints from cerebrospinal fluid (CSF), therefore actively promoting the contrast-to-noise ratio.
@@ 2 Theory
The proposed contrast-optimized basis was inspired by the concept of region-optimized virtual (ROVir) coils, where the generalized eigendecomposition was used to maximize the signal-to-interference ratio, showing promising results in suppressing unwanted signal in different MRI applications.9 We propose to use the generalized eigendecomposition to maximize the contrast-to-noise ratio between two types of tissue. This approach effectively rotates the SVD subspace, resulting in a contrast-optimized basis that promotes the contrast in the first and last coefficient images.
In the following, we will translate the ROVir formalism to subspace modeling. While this approach can be applied to any two types of tissue, we will outline the concept in the example of brain parenchyma and CSF and we will use simulated fingerprints to calculate the basis. First, we calculate an SVD and truncate it heuristically to the first 3 basis functions. This step ensures that the subspace covers most of the signal intensity, which minimizes artifacts from unmodeled signals.
SVD basis
proposed basis
Figure 1: Top: Basis functions $\mathbf{U}_{\text{SVD}}^{(1)}$ and $\mathbf{U}_{\text{opt}}^{(1)}$, which are used to reconstruct the low-resolution images for motion estimation. The original SVD basis maximizes the overall signal, resulting in a proton-density-like contrast (bottom left). The contrast-optimized basis was designed to maximize the signal arising from the brain parenchyma while minimizing the signal arising from the CSF, resulting in improved tissue contrast (bottom right).
@@
thus maximizing the signal fingerprints of the brain parenchyma while minimizing the signal fingerprints of the CSF. Since $\mathbf{C}_{\text{b}}$ and $\mathbf{C}_{\text{f}}$ are positive-semidefinite Hermitian-symmetric matrices, and $\mathbf{C}_{\text{f}}$ has full rank, there is a set of eigenvalues $\mathbf{\lambda}_{i}$ with linearly independent eigenvectors $\mathbf{w}_{i}$ such that
@@
Here, $i\in\{1,2,3\}$. Eq. (6) can be solved by calculating the generalized eigendecomposition 10 for $\mathbf{C}_{\text{b}}$ and $\mathbf{C}_{\text{f}}$ and order the generalized eigenvalues and eigenvectors so that $\mathbf{\lambda}_{1}\geq\mathbf{\lambda}_{2}\geq\mathbf{\lambda}_{3}$. Assuming normalized eigenvectors ($||\mathbf{w}_{i}||_{2}=1$), we can rotate the basis functions with
@@
The rotated subspace maximizes the parenchyma and minimizes the CSF signal in the first coefficient, while the last coefficient has the opposite properties. The Gram-Schmidt approach is then applied to ensure the orthogonality between the three bases, leaving the first basis function unchanged. The resulting brain basis function $\mathbf{U}_{\text{opt}}^{(1)}$ and coefficient image are shown in Fig. 1, which highlights the improved contrast between parenchyma and CSF compared to the SVD basis $\mathbf{U}_{\text{SVD}}^{(1)}$.
@@ 3.1 Data Acquisition
We acquired data from 11 healthy volunteers and 75 participants affected by mild Traumatic Brain Injury, for a total of 86 acquisitions. The participants were instructed to stay still during the scan. We used a 3 T Prisma scanner (Siemens Healthineers, Erlangen, Germany) on which we ran a 3D hybrid-state sequence11 optimized for quantitative magnetization transfer (qMT) imaging.12 Each RF pulse pattern is 4 s long and consists of an inversion pulse $\pi$, followed by a train of rectangular RF pulses with a varying flip angle and pulse duration and spaced 3.5 ms apart. We used six flip angle patterns, which were optimized to encode six biophysical MT parameters: $m_{0}^{s}$, $R_{1}^{f}$,$R_{2}^{f}$, $R_{\text{x}}$, $R_{1}^{s}$, and $T_{2}^{s}$.12 The sequence utilizes a 3D radial koosh-ball readout trajectory with reordered golden-angle increments 13, 14, 15, 16 and nominal resolution of 1 mm isotropic ($|k_{max}|=\pi/1\,mm$), sampling 1142 spokes per cycle of the RF pattern. Each RF pattern is repeated 30 times, for a total scan time duration of 12 min. For each subject, we also acquired a 3D MP-RAGE with 1 mm isotropic resolution. Informed consent was obtained prior to the scan in accordance with our Institutional Review Board.
Figure 2: Motion estimates derived from two different acquisitions: the “still” participant moved very little during the scan, whereas the other “moving” participant moved moderately. For the still participant, the motion scores were 0.47 mm and 0.5 mm with the SVD and proposed contrast-optimized basis, respectively, and for the moving participant 1.60 mm and 1.66 mm. Here, we show one representative translation (head-foot) and one representative rotation. All 6 motion parameters can be found in the Supporting Fig. S2.
Figure 3: Sagittal view of $m_{0}^{s}$ and $R_{x}$ maps of a scan with virtually no motion (a–f). The motion correction with both the SVD basis (b,e) and, in particular, the proposed basis (c,f) entailed little-to-no degradation of the apparent resolution compared to a reconstruction without motion correction (a,d). In the presence of motion (“moving” participant), the parameter maps reconstructed without motion correction are degraded (g,j). Motion correction with the original SVD basis substantially reduced motion artifacts (h,k), which is further improved with the proposed contrast-optimized basis (i,l).
@@ 3.2 Motion Estimation and Image Reconstruction
We aggregated all spokes from one 4 s RF cycle to reconstruct low-resolution coefficient images (4 mm isotropic) in the subspace of the SVD and the proposed contrast-optimized basis.17 A total variation (TV) penalty along time was used to mitigate undersampling artifacts and noise,15, 18 where the associated regularization strength was chosen heuristically based on the smoothness of the motion estimates and a visual inspection of the reconstructed low-resolution images. The reconstruction problem was solved using the Alternating Directions Method of Multipliers solver.19
The $N=180$ reconstructed low-resolution volumes were aligned separately for each flip angle pattern with the Statistical Parametric Mapping (SPM) software.20 More details regarding the extraction of the motion parameters can be found in Supporting Section S1. The resulting affine matrices were then used to rotate the k-space trajectory of each 4 s block. Translations were incorporated by multiplying the k-space data with a corresponding linear phase slope.
To compensate for the limitations of the 4 s temporal resolution of motion correction, we discarded blocks before and after large jumps in the motion estimates. To this end, we calculated a motion score between two time points $t$ and $\tau$ as:
@@
where $r$ is the spherical distance calculated on a sphere with radius $R=64\,mm$, $\theta_{t,\tau}$ is the angle of rotation extracted from the Euler angles of the estimated rotations, and $x_{t},y_{t},z_{t}$ and $x_{\tau},y_{\tau},z_{\tau}$ are the positions of the object at times $t$ and $\tau$21, 5 If the motion score was larger than a threshold, data from both neighboring time points were discarded in the final image reconstruction. The threshold was heuristically set at 0.75 mm, which visually resulted in the best image quality considering the trade-off between motion and undersampling artifacts. Supporting Fig. S1 provides an overview of the amount of data removed across the datasets.
Based on the motion-corrected k-space trajectory and data, the final image reconstruction was performed, also using subspace modeling.17, 22, 23 Here, we used a subspace optimized for the conservation of the Cramér–Rao bound,24 and we solved the reconstruction problem with the OptISTA algorithm.25 Finally, we estimated parameter maps with a neural network-based method.26, 27, 28
3.3 Data Analysis
We performed a region of interest (ROI) analysis to quantify the artifact level in the parameter maps. For our pulse sequence with a radial k-space readout, motion results, among others, in noise-like artifacts. Therefore, we used the standard deviation of each quantitative parameter in each ROI as a proxy for the artifact level.
We segmented the available MP-RAGE, and used Freesurfer to register them to the qMT maps, focusing on the following ROIs: global white matter, the pallidum, corpus callosum, and putamen. The ROI analysis was performed on 85/86 subjects, as the MP-RAGE of one dataset was non-diagnostic due to motion artifacts.
The qMT values for each ROI were then compared against the mean pair-wise motion score, which is defined as:

### v2

@@ Abstract
[Summary] Purpose: The long scan times of quantitative MRI techniques make them vulnerable to motion artifacts. For MR-Fingerprinting-like approaches, this problem can be addressed with self-navigated retrospective motion correction based on reconstructions in a singular value decomposition (SVD) subspace. However, the SVD promotes high signal intensity in all tissues, which limits the contrast between tissue types and ultimately reduces the accuracy of registration. The purpose of this paper is to rotate the subspace for maximum contrast between two types of tissue and improve the accuracy of motion estimates.
@@
Results: A comparative analysis shows that the contrast-optimized basis significantly improves the parenchyma-CSF contrast, leading to more accurate motion estimates and reduced artifacts in the quantitative maps.
@@
articletype: Submitted to Magnetic Resonance in Medicine00footnotetext: * E. Marchetto and S. Flassback contributed equally to this work.
@@ 1 Introduction
Magnetic resonance fingerprinting (MRF) is a multiparametric quantitative magnetic resonance imaging (qMRI) approach. Its key concept is a variation of sequence parameters between repetition times (TRs) to maintain a transient state of the magnetization. 1, 2
Motion-induced artifacts pose significant challenges in MRF-like experiments, whether due to physiological motion or involuntary motion due to pathological conditions. 3D acquisitions mitigate the risk of through-plane motion3—which is prohibitively difficult to correct retrospectively—but the associated long scan times make motion more likely.
Both real-time and retrospective motion correction strategies rely on navigators or external tracking devices to provide estimates of the subject’s motion.4 Fat-navigators have previously been proposed for robust retrospective motion correction in structural brain MRI,5, 6, 7 and they have successfully been implemented in a 3D MRF sequence.8 However, navigator-based motion correction techniques depend on the sequence timing, as they are typically acquired during waiting periods. Moreover, even fat-selective excitation pulses can perturb the spin dynamics via the magnetization transfer (MT) effect, which can compromise quantitative measurements.
On the other hand, external tracking devices are, in general, faster than navigator-based techniques and monitor the subject’s movements independently from the acquired sequence. Such systems include wearable sensors and optical tracking devices.9 However, despite being sequence agnostic, external tracking devices have seen limited use in the MRF community,10 likely due to the challenges associated with their clinical implementation, e.g., complexity of system setup.11
An alternative approach is self-navigated motion correction, which requires no additional data as the motion parameter estimation is derived from the acquired data itself. The Periodically Rotated Overlapping Parallel Lines with Enhanced Reconstruction (PROPELLER) is a widely used self-navigated imaging technique in which consecutively rotated blades are acquired around the k-space center.12 Each blade can be reconstructed into a low-resolution image, which is then used to detect and correct motion artifacts. PROPELLER-like acquisition schemes require long acquisition times and are not well suited for dynamic imaging protocols in which flip angles and pulse durations change with every TR.
Self-encoded FID navigators have also shown compelling results for self-navigated motion correction using a 3D radial trajectory.13 However, they necessitate the acquisition of prescan normalized data prior to the scan to calibrate the motion model. Additionally, the associated spin perturbations would likely need to be accounted for in the model when combining FID navigators with quantitative MRI.
In 3D MRF acquisitions, a self-navigated motion correction approach was previously proposed. 14 Kurzawski et al. reconstructed brain navigators from 7 s segments in a sub-space spanned by the truncated singular value decomposition (SVD) subspace. This method leverages the low-rank nature of the underlying data to produce coefficient images, which are then used to extract the motion estimates utilizing rigid registration. However, the SVD promotes high signal intensity in all types of tissue, which limits the contrast and can ultimately reduce the accuracy of the registration and extracted motion estimates.
In this work, we aim to improve the accuracy of motion estimates by deriving a basis that maximizes the signal of the fingerprints corresponding to brain parenchyma (i.e., white and gray matter), while minimizing the signal of fingerprints from cerebrospinal fluid (CSF), therefore actively promoting the contrast-to-noise ratio.
@@ 2 Theory
The proposed contrast-optimized basis was inspired by the concept of region-optimized virtual (ROVir) coils, where the generalized eigendecomposition was used to maximize the signal-to-interference ratio, yielding promising results in suppressing unwanted signals in various MRI applications.15 We propose to use the generalized eigendecomposition to maximize the contrast-to-noise ratio between two types of tissue. This approach effectively rotates the SVD subspace, resulting in a contrast-optimized basis that promotes the contrast in the first and last coefficient images.
In the following, we will translate the ROVir formalism to subspace modeling. While this approach can be applied to any two types of tissue, we will outline the concept in the example of brain parenchyma and CSF, and we will use simulated fingerprints to calculate the basis. First, we calculate an SVD and truncate it heuristically to the first 3 basis functions. This step ensures that the subspace encompasses most of the signal intensity, which minimizes artifacts from unmodeled signals.
Figure 1: Top: Exemplary basis functions $\mathbf{U}_{\text{SVD}}^{(1)}$ and $\mathbf{U}_{\text{opt}}^{(1)}$, which are used to reconstruct low-resolution images for motion estimation (bottom), derived by aggregating the radial spokes acquired during one 4 s RF cycle (cf. Sec. 3.2). The SVD basis maximizes the signal across all tissues, discouraging contrast between them. In contrast, the proposed basis directly maximizes the contrast between tissues (in this case, brain parenchyma and CSF). We note that the contrast in the SVD basis exhibits considerable variability between sequences, choices of parameters used for simulating each tissue, etc., since contrast is not part of the optimization objective. Here, we show a case with particularly low contrast, more examples can be found in Supplementary Fig. S1.
@@
thus, maximizing the signal fingerprints of the brain parenchyma while minimizing the signal fingerprints of the CSF. Since $\mathbf{C}_{\text{b}}$ and $\mathbf{C}_{\text{f}}$ are positive-semidefinite Hermitian-symmetric matrices, and $\mathbf{C}_{\text{f}}$ has full rank, there is a set of eigenvalues $\mathbf{\lambda}_{i}$ with linearly independent eigenvectors $\mathbf{w}_{i}$ such that
@@
Here, $i\in\{1,2,3\}$. Eq. (6) can be solved by calculating the generalized eigendecomposition 16 for $\mathbf{C}_{\text{b}}$ and $\mathbf{C}_{\text{f}}$ and order the generalized eigenvalues and eigenvectors so that $\mathbf{\lambda}_{1}\geq\mathbf{\lambda}_{2}\geq\mathbf{\lambda}_{3}$. Assuming normalized eigenvectors ($||\mathbf{w}_{i}||_{2}=1$), we can rotate the basis functions with
@@
The subspace maximizes the parenchyma and minimizes the CSF signal in the first coefficient, while the last coefficient has the opposite properties. The Gram-Schmidt approach is then applied to ensure the orthogonality between the three bases, leaving the first basis function unchanged. The resulting brain basis function $\mathbf{U}_{\text{opt}}^{(1)}$ and coefficient image are shown in Fig. 1, which highlights the improved contrast between parenchyma and CSF compared to the SVD basis $\mathbf{U}_{\text{SVD}}^{(1)}$.
@@ 3.1 Data Acquisition
The acquired data comprised 11 healthy control scans and 75 scans from participants affected by mild Traumatic Brain Injury, for a total of 86 acquisitions. The participants were instructed to stay still during the scan. We used a 3 T Prisma scanner (Siemens Healthineers, Erlangen, Germany) on which we ran a 3D hybrid-state sequence17 optimized for quantitative magnetization transfer (qMT) imaging.18 Each RF pulse pattern is 4 s long and consists of an inversion pulse $\pi$, followed by a train of 1141 rectangular RF pulses with a varying flip-angle and pulse duration and spaced 3.5 ms apart.
Based on Henkelman’s two-pools model,19 the free-water pool (denoted by the superscript $f$) exchanges magnetization with the semi-solid spin pool (denoted by the superscript $s$) at the rate $R_{x}$. Each pool has a fractional pool size, denoted as $m_{0}^{f}$ and $m_{0}^{s}$, where $m_{0}^{f}+m_{0}^{s}=1$. Additionally, each pool’s spin dynamics is captured by respective longitudinal ($R_{1}^{f}$, $R_{1}^{s}$) and transverse ($R_{2}^{f}$, $T_{2}^{s}$) relaxation rates, which are the inverse of respective relaxation times. We used six different flip-angle patterns, optimized to encode these six core biophysical MT parameters: $m_{0}^{s}$, $R_{1}^{f}$, $R_{2}^{f}$, $R_{\text{x}}$, $R_{1}^{s}$, and $T_{2}^{s}$ (note that we use the relaxation time $T_{2}^{s}$ for historic reasons).18
The sequence utilizes a 3D radial koosh-ball readout trajectory with reordered golden-angle increments 20, 21, 22, 23 and nominal resolution of 1 mm isotropic ($|k_{max}|=\pi/1\,mm$). Each RF pattern is repeated 30 times, for a total scan time duration of 12 min. For each subject, we also acquired a 3D MP-RAGE with 1 mm isotropic resolution. Informed consent was obtained prior to the scan in accordance with our Institutional Review Board.
Figure 2: Estimates of simulated motion, using the SVD and proposed basis. The ground truth motion was obtained from an online database and has a motion score of 1.77 mm. The magnifications highlight improvements of the proposed over the SVD basis, which is corroborated by a reduction in the RMSE score from 0.38 mm to 0.35 mm (lower bound of 0.3 mm calculated between the ground truth and the downsampled ground truth). Here, we show one representative translation (head-foot) and one representative rotation. All 6 motion parameters can be found in the Supporting Fig. S4.
@@ 3.2 Motion Estimation and Image Reconstruction
We aggregated all acquired spokes from each 4 s RF cycle to reconstruct low-resolution coefficient images (4 mm isotropic) in the subspace of the SVD and the proposed contrast-optimized basis.24 The reconstruction problem can be formulated as:
$\hat{x}=\text{argmin}_{x}\frac{1}{2}\mid\mid{Ex-y}\mid\mid^{2}_{2}+\lambda_{t}\mid\mid{\nabla_{t}x}\mid\mid_{1}$ (8)
where $E$ represents the encoding operator, which combines gridding, Fourier transform, and coil sensitivities. We used a total variation (TV) penalty along time to mitigate undersampling artifacts and noise,22, 25 where the associated regularization strength $\lambda$ was chosen based on the simulations described in Sec. 3.3.4. The finite difference operator is denoted by $\nabla_{t}$. The reconstruction problem was solved using the Alternating Directions Method of Multipliers solver.26 Examples of low-resolution coefficient images reconstructed either using the SVD basis or the proposed contrast-optimized basis for each of the six flip-angle patterns are shown in Fig. 1 and Supporting Fig. S1.
The $N=180$ reconstructed low-resolution volumes were rigidly co-registered separately for each flip-angle pattern with the Statistical Parametric Mapping (SPM) software.27 More details regarding the extraction of the motion parameters can be found in Supporting Sec. S1. The resulting affine matrices were then used to rotate the k-space trajectory of each 4 s block. Translations were incorporated by multiplying the k-space data with a corresponding linear phase slope.
Due to the limited 4 s temporal resolution of the motion correction, which assumes no motion occurs within that time frame, we discarded the 4 s blocks immediately before and after large jumps in the motion estimates. To this end, we calculated a motion score between two time points $t$ and $\tau$ as:
@@
where $r$ is the spherical distance calculated on a sphere with radius $R=64\,mm$, $\theta_{t,\tau}$ is the angle of rotation extracted from the Euler angles of the estimated rotations, and $x_{t},y_{t},z_{t}$ and $x_{\tau},y_{\tau},z_{\tau}$ are the positions of the object at times $t$ and $\tau$28, 6 If the motion score was larger than a threshold, data from both neighboring time points were discarded in the final image reconstruction. The threshold was heuristically set at 1.5 mm, which visually resulted in the best image quality considering the trade-off between motion and undersampling artifacts. Supporting Fig. S2 provides an overview of the amount of data removed across the datasets.
Based on the motion-corrected k-space trajectory and data, the final image reconstruction was performed, also using subspace modeling.24, 29, 30 Here, we used a subspace optimized for the conservation of the Cramér–Rao bound,31 and we solved the reconstruction problem, which includes a locally low-rank penalty, with the OptISTA algorithm.32 Finally, we estimated parametric maps with a neural network-based method.33, 34, 35
3.3 Motion Simulations
To evaluate the proposed approach, we performed motion simulations by corrupting a reference dataset and performing motion correction on the corrupted dataset. By comparing motion estimates with the ground truth motion, we evaluated the accuracy of motion estimates.
3.3.1 Reference Dataset
As reference dataset, we selected the participant with the least amount of motion ($\bar{M}=0.22$ mm) from our pool of 86 acquisitions, where
@@ 3.3.1 Reference Dataset
Figure 3: Sagittal view of $m_{0}^{s}$ and $R_{x}$ maps of the selected reference image (no motion added; Sec. 3.3.1) and with simulated motion (b, f), which show strong degradations. In contrast, both the SVD basis (c, g) and the proposed contrast-optimized basis (d, h) significantly improve image quality, with the latter visually agreeing most closely with the reference maps (a, e). The RMSE for each parameter map, compared to the reference (no motion added), is shown in the top left of the respective map. The RMSE was calculated for the entire 3D brain volume, defined by the brain mask. The remaining parametric maps are shown in Supporting Fig. S5.
3.3.2 Ground Truth Motion
Ground truth motion parameters were sourced from a publicly available database (http://34.88.15.16/webapps/home/session.html?app=BrainMRIMotionDB), initiated by the motion correction research community. We selected seven motion datasets with unintended head motion that were acquired from pediatric patients. The motion parameters were recorded at a sampling rate of 30 Hz using an optical tracking system, and we interpolated them to our sampling interval of $T_{\text{R}}=3.5\;\text{ms}$ using linear B-splines. We corrupted the reference dataset with each of the seven sets of motion parameters by rotating each radial k-space spoke as specified by the rotation parameters and by adding a linear phase slope to the corresponding data as specified by the translation parameters.
3.3.3 Motion Estimation
We estimated the motion parameters as described in Sec. 3.2. Although we selected the reference image based on its low motion score, it is not entirely motion-free. To account for the baseline motion, we estimated the motion parameters from the uncorrupted and the corrupted reference data. We computed the difference by dividing the affine matrices, and we compared this difference to the ground truth motion parameters. This pipeline was performed for each of the seven sets of ground truth motion parameters and separately for the SVD and the proposed contrast-optimized basis.
Figure 4: Motion estimates of two extreme acquisitions: the scan with the $0^{th}$ percentile motion score (0.38 mm and 0.41 mm with the SVD and contrast-optimized basis; no data removed), and the scan with the $100^{th}$ percentile motion score (10.32 mm and 10.53 mm; 16.11% and 12.22% of data removed respectively). Here, we show one representative translation (head-foot) and one representative rotation. All 6 motion parameters can be found in the Supporting Fig. S2.
3.3.4 $\lambda$ Optimization
The central tuning parameter in our pipeline is the regularization parameter $\lambda$ for the low-resolution reconstruction (see Eq. 8). We used the simulated motion for an objective selection of $\lambda$. As the figure of merit, we defined the RMSE score as the $\ell_{2}$-norm of the root mean squared error (RMSE) between the estimated and ground truth motion parameters:
$\lambda_{\text{opt}}=\arg\min_{\lambda}\sqrt{\sum_{i=1}^{3}\text{MSE}_{\text{trans},i}(\lambda)+R^{2}\sum_{i=1}^{3}\text{MSE}_{\text{rot},i}(\lambda)}.$ (13)
Here, $i$ loops over the three translation and rotation parameters, and the radius $R=64\,\text{mm}$ converts rotations to translations (cf. Eqs. (9)—(11)). We determined an optimal $\lambda$ for each of the six flip-angle patterns separately using three of the seven ground truth motion datasets, testing $\lambda\in\{ke^{-4}\mid k=1,\dots,10\}$, and choosing the median $\lambda$ between the three datasets (see Supporting Fig. S3)
Finally, we evaluated the motion estimation accuracy of the SVD basis and the proposed contrast optimized basis by comparing the RMSE score across the remaining four motion patterns.
3.4 ROI Analysis
To evaluate the performance of the motion correction across our 86 datasets, we performed a region of interest (ROI) analysis. As our pulse sequence uses a radial k-space readout, motion primarily introduces noise-like artifacts. Therefore, we calculated the standard deviation of each quantitative parameter per ROI as a proxy for the artifact level.
We registered the available MP-RAGE to the qMT maps using Freesurfer, and segmented the former, focusing on the following ROIs: global white matter, pallidum, corpus callosum, and putamen. The ROI analysis was performed on 85/86 subjects, as one MP-RAGE was excluded due to failed segmentation caused by extensive motion artifacts.
We compared the standard deviation of each qMT parameter and each ROI to the mean pair-wise motion score (cf. Sec. 3.3.1). We performed linear regression to determine whether the artifact level increases with motion, hypothesizing that the corresponding slope is reduced when motion correction is performed with the proposed basis.
Figure 5: Sagittal view of $m_{0}^{s}$ and $R_{x}$ maps of a scan with minimal motion (a–f). The motion correction with the SVD basis (b,e) introduced subtle artifactual hyperintensities in the $R_{x}$ map (magnification in e), which are reduced when using the proposed basis (c,f). In the presence of strong motion (100th percentile motion score), the parametric maps reconstructed without motion correction are severely degraded (g,j). Motion correction with the original SVD basis substantially reduced motion artifacts (h,k), which is further improved with the proposed contrast-optimized basis (i,l). Like in Fig. 3a, the $m_{0}^{s}$ map is expected to have an MP-RAGE-like contrast, while the exchange rate $R_{x}$ is expected to be near-isointense between gray and white matter, like in Fig. 3e.18